Solve the following differential equations : (x + cosy)dx + (-xsiny)=0


1
Expert's answer
2022-06-03T13:25:20-0400
"\dfrac{\partial Q}{\partial x}=-\sin y=\dfrac{\partial P}{\partial y}"

"u=\int (x+\cos y)dx+\varphi(y)"

"=\dfrac{x^2}{2}+x\cos y+\varphi(y)"

"\dfrac{\partial u}{\partial y}=-x\sin y+\varphi '(y)=-x\sin y"

"\varphi '(y)=0"

"\varphi(y)=-C"


"u=\dfrac{x^2}{2}+x\cos y-C"

The general solution of the exact differential equation is given by


"\dfrac{x^2}{2}+x\cos y=C"